A wave is modeled with the function y(x,t) = (0.58m) cos((0.21rad/m)x − (0.03rad/s)t + π/3). Find the wave speed.

Round your answer to 2 decimal places


1
Expert's answer
2021-04-26T17:03:49-0400
"v=\frac{\omega}{k}\\v=\frac{0.03}{0.21}=0.14\frac{m}{s}"